OpenStudy (anonymous):

A block, initially at rest on a friction-free horizontal plane, is acted on by a horizontal force of 25 N. If during 4.0 seconds the block travels 40.0 m, what is its mass?

OpenStudy (abhisar):

Hello @dray71837 , Welcome to Openstudy !

OpenStudy (abhisar):

Using the equation, \(\sf s=ut+\frac{1}{2}at^2\) where, s=distance u=initial velocity a=acceleration and t=time We can calculate the value of acceleration as follows. 40=0.5*a*16 =>a=5 m/\(\sf s^2\) \(\sf \because\) Force=acceleration*Mass \(\sf \therefore Mass=\frac{Force}{acceleration}=\frac{25}{5}=5Kg\)

OpenStudy (anonymous):

Thank you so much! @Abhisar